OpenStudy (anonymous):

Suppose that 4 ≤ f '(x) ≤ 5 for all values of x. What are the minimum and maximum possible values of f(3) − f(1)? ?≤ f(3) − f(1) ≤? Im not really sure how to do this, the examples that i have aren't like this one, please help!

OpenStudy (anonymous):

4 ≤ f '(x) ≤ 5 means the slope of the tangent line is between 4 and 5. The slope of the secant line from 1 to 3 is \[m = \frac{ f(3)-f(1) }{ 3-1 }=\frac{ f(3)-f(1) }{ 2 }\] So I think to get your answers you need to solve \[4 ≤\frac{ f(3)-f(1) }{ 2 }≤5\]

OpenStudy (anonymous):

what would i plug in for f(3) though?

OpenStudy (freckles):

From your question it looks like you want to solve for f(3)-f(1)

OpenStudy (freckles):

do you know how to undo division an equation... hint multiply by something that un does it

OpenStudy (freckles):

and this an inequality not an equation but basically works the same way

OpenStudy (anonymous):

okay so i get 8<f(3)-f(1)<10

OpenStudy (anonymous):

got it thanks!!